OpenStudy (anonymous):

square root 10 x square root 6

OpenStudy (anonymous):



OpenStudy (anonymous):

Thats what calclators are for!!! lol

OpenStudy (anonymous):

Ok what do you do when you multiply square roots?

OpenStudy (anonymous):

You simplify them by factoring them @kellyking

OpenStudy (anonymous):

you simplify them but i dont know how

OpenStudy (anonymous):

So do that..... you will get your answer

OpenStudy (anonymous):

but arent 10 and 6 already simplified

OpenStudy (anonymous):

Ok well I will tell you! Combine them!

OpenStudy (anonymous):

So it would be the square root of 10*6

OpenStudy (zzr0ck3r):

\(\sqrt{10}*\sqrt{6}=\sqrt{60}=\sqrt{4*15}=\sqrt{4}\sqrt{15}=2\sqrt{15}\)

OpenStudy (anonymous):

Dont you just use 5 and 2 to simplify 10 and you can use 3 and 2 to simplify 6

OpenStudy (anonymous):

Yes but i was only saying the first step

OpenStudy (anonymous):

oh..sorry dude

OpenStudy (anonymous):

And there are so many ways to factor it

OpenStudy (anonymous):

I gtg bye guys!

OpenStudy (zzr0ck3r):

\(\sqrt{ab}=\sqrt{a}\sqrt{b}\\\sqrt{\frac{a}{b}}=\frac{\sqrt{a}}{\sqrt{b}}\)

OpenStudy (anonymous):

just simplfy and then multiply them @kellyking

OpenStudy (anonymous):

Hehe I was right!!!

OpenStudy (anonymous):

thanks guys!!